OpenStudy (rachelallen2k16):

f(x) = two divided by quantity x squared minus two x minus three find the domain and range x and y intercepts horizontal asymptotes vertical asymptotes

OpenStudy (mathmale):

Rachel, could you possibly share this function f(x) written in math symbols instead of in English? It's all too easy to misinterpret functions expressed in words. If you must use words, then use parentheses to indicate which groups of words go together.

OpenStudy (rachelallen2k16):

f(x)=2/x^2-2x-3

OpenStudy (sshayer):

find the value of x where denominator not zero.

OpenStudy (mathmale):

Rachel, Your f(x)=2/x^2-2x-3 is much easier to read. Thank you. Just one more request of you: Please enclose the denominator in parentheses. Otherwise some readers may assume yuo're dividing 2 by x^2 and then subtracting 2x+3 from the result. \[f(x)=2/x^2-2x-3\rightarrow \frac{ 2 }{ (x^2-2x-3) }=f(x)\]

OpenStudy (mathmale):

Domain: The set of input (x-) values for which the function is defined. Your function is not defined if the denom. happens to be zero. Thus, you must set the denom. = 0 and solve the resulting quadratic equation for x. You should get two answers. What will you do with them, so as to convey your domain clearly and effectively?